# HodgeFormer: Transformers for Learnable Operators on Triangular Meshes through Data-Driven Hodge Matrices

### v3

@@ 5.2 Experiments and Results
5.3 Model Variations Study
To evaluate the importance of different architecture components and parameter choices, we varied our base model in different ways and measured its performance on a segmentation task. Specifically, in Tab. 5, we compare the main architecture layers HodgeFormer and a vanilla Transformer layer along with different embedding layers, i.e., the feed-forward and the neighbor embedding layers.
In Tab. 4, we test combinations of different input features based on their type by distinguishing between coordinates, normals, and areas. In Table  6, we evaluate the effect of the number of neighbors taken into account by the sparse attention component on the performance of the HodgeFormer layers.
For all experiments, we consider an architecture of 4 layers acting only on mesh vertices with embedding dimension 256 and hidden dimension 512, and a learning rate of 5e-4, and we report test accuracy results, averaged over 5 experiment trials with random training splits. We also consider $\sqrt{n}$ neighbors for each set of $n$ mesh elements with the exception of the experiment of Tab. 6.
In Tab. 5, the HodgeFormer layers clearly outperform vanilla Transformer layers. At the same time, embedding input features while considering their neighbors offers an edge in initial training steps and results in an overall better performance. In Tab. 6, we observe the effect of the number of neighbors for vertices and edges on the HodgeFormer performance. For $s_{v}=1,s_{e}=1$, the performance drops significantly as the attention mechanism takes into account only self-contributions for vertices and edges. For $s_{v}=32,s_{e}=48$, which approximates $\sqrt{n}$ for meshes in the COSEG vases dataset, the performance is high. Yet, for larger numbers of neighbors, we get diminishing returns in model performance. In Tab. 4, we observe that the HodgeFormer layer is capable of achieving satisfactory results using simple coordinate features, although exhibiting lower performance and slower convergence. However, using normals as the sole feature set proves insufficient, resulting in models with poor performance and limited generalization capability.
@@ 5.4 Efficiency
The HodgeFormer architecture uses standard linear algebra operations throughout preprocessing, training, and inference. Preprocessing involves data loading, augmentation, and BFS operations for local neighborhood extraction (enabling sparse attention), all performed on-the-fly during data loading without separate precomputation, supporting continuous data streaming across multiple cores.
Tab. 7 compares the runtime performance of HodgeFormer with other recent methods on the human segmentation task, under typical training and testing configurations. HodgeFormer has an attractive runtime profile compared to other methods with fast training, high GPU utilization and no precomputation steps. It should be noted that a significant part of its execution time is spent in I/O and preprocessing, mostly in the BFS-based neighborhood extraction.
@@
Transformer Embedding Vases
Vanilla Transformer FNN 84.82%
Vanilla Transformer Neighbors + FNN 87.28%
HodgeFormer FNN 91.67%
HodgeFormer Neighbors + FNN 92.02%
@@
MeshCNN 1.36 1.09 60.40 0.19
Laplacian2Mesh 2.96 0.13 7.98 0.95
@@ 6 Discussion, limitations and conclusion
In this work, we draw connections between discrete Hodge operators and Transformer-style attention mechanisms and propose a general architecture for deep learning on triangular meshes named HodgeFormer. Our architecture omits the construction of commonly used techniques such as expensive spectral features, eigenvalue decomposition operations, and complex complementary structures. Despite these omissions, HodgeFormer retains its expressiveness while fully leveraging the hardware-friendly properties of the Transformer architecture. In addition, by employing sparse attention mechanisms, the HodgeFormer architecture results in an overall complexity of $O(n^{1.5}d)$. HodgeFormer was tested on classification and segmentation tasks, demonstrating performance comparable to state-of-the-art methods. Strategies such as patching in MeshMAE [21] or mesh simplification preprocessing in MeshWalker [16] are part of future research on large-scale meshes and large-scale unsupervised training. Additionally, while this research focused on Hodge Laplacian operators, our framework can be extended to other families of discrete operators, such as the Dirac operator. Finally, compared to a standard Transformer, our architecture has higher memory requirements due to the use of multiple attention mechanisms per layer. This requirement is partially mitigated by the employment of sparse attention. Although the proposed architecture has low complexity and is fully parallelizable, its implementation relies heavily on gather operations. Investigation of preprocessing strategies for input meshes with reordering algorithms such as the Cuthill-McKee [4] may allow the application of contiguous operations and enable better hardware utilization.
@@ A.1 Multi-head Hodge Attention on vertices, edges and faces
This section presents a detailed analysis of the HodgeFormer attention mechanism, demonstrating how the multi-head Hodge Attention component operates across different mesh elements. By combining the Hodge star operator formulation (Eq. 12) with the attention-based operators (Eq. 6), the Hodge Laplacians $L_{v}$, $L_{e}$, and $L_{f}$ for vertices (0-forms), edges (1-forms), and faces (2-forms) are reformulated as the attention-based expressions given in equations 14, A.1, and 16, respectively.
$L_{v}\enspace:=\enspace$ $\,\star_{0}^{-1}(x_{v})\cdot d_{0}^{T}\cdot\star_{1}(x_{e})\cdot d_{0}$
$:=\enspace$ $\sigma\left({\frac{Q_{v}K_{v}^{T}}{\sqrt{d_{k}}}}\right)\cdot d_{0}^{T}\cdot\sigma\left({\frac{Q_{e}K_{e}^{T}}{\sqrt{d_{k}}}}\right)$ (14)
$L_{e}\enspace:=\enspace$ $\,d_{0}\cdot\star_{0}^{-1}(x_{v})\cdot d_{0}^{T}\cdot\star_{1}(x_{e})\,+$
@@
$:=\enspace$ $\,d_{0}\cdot\sigma\left({\frac{Q_{v}K_{v}^{T}}{\sqrt{d_{k}}}}\right)\cdot d_{0}^{T}\cdot\sigma\left({\frac{Q_{e1}K_{e1}^{T}}{\sqrt{d_{k}}}}\right)\,+$
$\quad\,\,\,\,\sigma\left({\frac{Q_{e2}K_{e2}^{T}}{\sqrt{d_{k}}}}\right)\cdot d_{1}^{T}\cdot\sigma\left({\frac{Q_{f}K_{f}^{T}}{\sqrt{d_{k}}}}\right)\cdot d_{1}$ (15)
$L_{f}\enspace:=\enspace$ $\,d_{1}\cdot\star_{1}^{-1}(x_{e})\cdot d_{1}^{T}\cdot\star_{2}(x_{f})$
$:=\enspace$ $\,d_{1}\cdot\sigma\left({\frac{Q_{e2}K_{e2}^{T}}{\sqrt{d_{k}}}}\right)\cdot d_{1}^{T}\cdot\sigma\left({\frac{Q_{f}K_{f}^{T}}{\sqrt{d_{k}}}}\right)$ (16)
@@
$x_{v}$ $=L_{v}\cdot V_{v}\quad\text{(vertex features)}$ (17)
$x_{e}$ $=L_{e}\cdot V_{e}\quad\text{(edge features)}$ (18)
$x_{f}$ $=L_{f}\cdot V_{f}\quad\text{(face features)}$ (19)
This formulation ensures that the attention mechanism respects the topological structure of the mesh while enabling information flow between different dimensional elements (vertices, edges, faces). The formation of these updated features is illustrated in Figures 7, 8, and 6, which visualize the computational flow described in equations (17), (18), and (19), respectively.
Figure 6: Multi-head Hodge Attention applied to latent face features $x_{f}$. The multi-head attention mechanism learns data-driven Hodge Star matrices $\star_{1}^{-1}$ and $\star_{2}$.
Figure 7: Multi-head Hodge Attention applied to latent vertex features $x_{v}$. The multi-head attention mechanism learns data-driven Hodge Star matrices $\star_{0}^{-1}$ and $\star_{1}$.
Figure 8: Multi-head Hodge Attention applied to latent edge features $x_{e}$. The multi-head attention mechanism learns data-driven Hodge Star matrices $\star_{0}^{-1}$ and $\star_{1}$.
@@ A.2 Computational complexity and memory requirements
The resulting complexity is $O(nd^{2})+O(n^{1.5}d)$ with $d$ fixed and usually much smaller than $n$. In comparison, the eigendecomposition of a sparse $(n,n)$ matrix has complexity $O(kn^{2})$ for calculating the first $k$ eigenvectors. This holds for Laplacian positional embeddings as well as for spectral features such as Heat Kernel Signature (HKS). Tab. 8 presents experiments evaluating the computational and memory requirements.
@@
Table 8: Metrics for different mesh sizes with respect to the number of vertices, for a 1-layer HodgeFormer (HF) end-to-end architecture (training and inference) as well as a standalone HodgeFormer layer (inference). The model operates on vertices with a latent embedding dimension $d=256$ and hidden MLP dimension of $d_{h}=512$. Measured on an Nvidia RTX 4090 GPU.
Layers Vases
@@
Table 9: Variation experiments comparing different strategies of mixing HodgeFormer with Transformer layers and their effect on the performance, evaluated on the COSEG Vases dataset.
@@
Appendix B Results
B.1 Ablation study on mixing strategies
This section examines the effect of mixing HodgeFormer and Transformer layers on segmentation performance. Table 9 summarizes the performance on the COSEG Vases dataset under varying layer compositions. We denote $N_{H}$ as the number of HodgeFormer layers and $R_{H:T}=(N_{H}:N_{T})$ the ratio of the number of HodgeFormer layers to Vanilla Transformer layers. As shown in Table 9, increasing the number of Transformer layers while keeping the number of HodgeFormer layers fixed ($N_{H}=4$) shows a general improvement, with the best result (93.04%) obtained when mixing six layers in total ($R_{H:T}=4:2$).
B.2 Eigenfunctions of the corresponding learned operator
We compute the eigenfunctions of the corresponding learned operators and visualize them on the mesh geometry, providing valuable insights into how the HodgeFormer captures and processes geometric features at different scales and orientations. Figure 10 presents the first five eigenfunctions of a COSEG Aliens dataset sample. Figure 11 presents the learned attention maps for two sample meshes of COSEG Aliens dataset measured in layer 6 of the architecture.
Figure 10: First few eigenfunctions of the corresponding learned operator.
Figure 11: Learned attention maps from two samples of the COSEG Aliens dataset

### v4

@@ 5.4 Efficiency
For the HodgeFormer architecture, preprocessing involves data loading, augmentation, and BFS operations for local neighborhood extraction (enabling sparse attention), all performed on-the-fly during data loading without separate precomputation, supporting continuous data streaming across multiple cores. Tab. 7 compares the runtime performance of HodgeFormer with other recent methods on the human segmentation task, under typical training and testing configurations. HodgeFormer has an attractive runtime profile compared to other methods with fast training, high GPU utilization and no precomputation steps, with a significant part of its execution time spent in I/O and preprocessing, mostly in the BFS-based neighborhood extraction.
Transformer Embedding Vases
Vanilla Transformer FNN 84.82%
Vanilla Transformer Neighbors + FNN 87.28%
HodgeFormer FNN 91.67%
HodgeFormer Neighbors + FNN 92.02%
@@
MeshCNN [11] 1.36 1.09 60.40 0.19
Laplacian2Mesh[18] 2.96 0.13 7.98 0.95
@@ 6 Discussion, limitations and conclusion
In this work, we draw connections between discrete Hodge operators and Transformer-style attention mechanisms and propose a general architecture for deep learning on triangular meshes named HodgeFormer. Our architecture omits the construction of commonly used techniques such as expensive spectral features, eigenvalue decomposition operations, and complex complementary structures. Despite these omissions, HodgeFormer retains its expressiveness while fully leveraging the hardware-friendly properties of the Transformer architecture. In addition, by employing sparse attention mechanisms, the HodgeFormer architecture results in an overall complexity of $O(n^{1.5}d)$. HodgeFormer was tested on classification and segmentation tasks, demonstrating performance comparable to state-of-the-art methods. Strategies such as patching in MeshMAE [21] or mesh simplification preprocessing in MeshWalker [16] are part of future research on large-scale meshes and large-scale unsupervised training. Compared to a standard Transformer, our architecture has higher memory requirements due to the use of multiple attention mechanisms per layer. This requirement is partially mitigated by the employment of sparse attention. Although the proposed architecture has low complexity and is fully parallelizable, its implementation relies heavily on gather operations. Investigation of preprocessing strategies for input meshes with reordering algorithms such as the Cuthill-McKee [4] may allow the use of contiguous operations and enable better hardware utilization.
@@ A.1 Multi-head Hodge Attention on vertices, edges and faces
This section presents a detailed analysis of the HodgeFormer attention mechanism, demonstrating how the multi-head Hodge Attention component operates across different mesh elements. By combining the Hodge star operator formulation (Eq. 12) with the attention-based operators (Eq. 6), the Hodge Laplacians $L_{v}$, $L_{e}$, and $L_{f}$ for vertices (0-forms), edges (1-forms), and faces (2-forms) are reformulated as the attention-based expressions given in equations S1, A.1, and S3, respectively.
$L_{v}\kern 5.0pt:=\kern 5.0pt$ $\,\star_{0}^{-1}(x_{v})\cdot d_{0}^{T}\cdot\star_{1}(x_{e})\cdot d_{0}$
$:=\kern 5.0pt$ $\sigma\left({\frac{Q_{v}K_{v}^{T}}{\sqrt{d_{h}}}}\right)\cdot d_{0}^{T}\cdot\sigma\left({\frac{Q_{e}K_{e}^{T}}{\sqrt{d_{h}}}}\right)$ (S1)
$L_{e}\kern 5.0pt:=\kern 5.0pt$ $\,d_{0}\cdot\star_{0}^{-1}(x_{v})\cdot d_{0}^{T}\cdot\star_{1}(x_{e})\,+$
@@
$:=\kern 5.0pt$ $\,d_{0}\cdot\sigma\left({\frac{Q_{v}K_{v}^{T}}{\sqrt{d_{h}}}}\right)\cdot d_{0}^{T}\cdot\sigma\left({\frac{Q_{e1}K_{e1}^{T}}{\sqrt{d_{h}}}}\right)\,+$
$\quad\,\,\,\,\sigma\left({\frac{Q_{e2}K_{e2}^{T}}{\sqrt{d_{h}}}}\right)\cdot d_{1}^{T}\cdot\sigma\left({\frac{Q_{f}K_{f}^{T}}{\sqrt{d_{h}}}}\right)\cdot d_{1}$ (S2)
$L_{f}\kern 5.0pt:=\kern 5.0pt$ $\,d_{1}\cdot\star_{1}^{-1}(x_{e})\cdot d_{1}^{T}\cdot\star_{2}(x_{f})$
$:=\kern 5.0pt$ $\,d_{1}\cdot\sigma\left({\frac{Q_{e2}K_{e2}^{T}}{\sqrt{d_{h}}}}\right)\cdot d_{1}^{T}\cdot\sigma\left({\frac{Q_{f}K_{f}^{T}}{\sqrt{d_{h}}}}\right)$ (S3)
@@
$x_{v}$ $=L_{v}\cdot V_{v}\quad\text{(vertex features)}$ (S4)
$x_{e}$ $=L_{e}\cdot V_{e}\quad\text{(edge features)}$ (S5)
$x_{f}$ $=L_{f}\cdot V_{f}\quad\text{(face features)}$ (S6)
Figure S1: Multi-head Hodge Attention applied to latent face features $x_{f}$. The multi-head attention mechanism learns data-driven Hodge Star matrices $\star_{1}^{-1}$ and $\star_{2}$.
Figure S2: Multi-head Hodge Attention applied to latent vertex features $x_{v}$. The multi-head attention mechanism learns data-driven Hodge Star matrices $\star_{0}^{-1}$ and $\star_{1}$.
Figure S3: Multi-head Hodge Attention applied to latent edge features $x_{e}$. The multi-head attention mechanism learns data-driven Hodge Star matrices $\star_{0}^{-1}$ and $\star_{1}$.
This formulation ensures that the attention mechanism respects the topological structure of the mesh while enabling information flow between different dimensional elements (vertices, edges, faces). The formation of these updated features is illustrated in Figures S2, S3, and S1, which visualize the computational flow described in equations (S4), (S5), and (S6), respectively.
@@ A.2 Computational complexity and memory requirements
The resulting complexity is $O(nd^{2})+O(n^{1.5}d)$ with $d$ fixed and usually much smaller than $n$. In comparison, the eigendecomposition of a sparse $(n,n)$ matrix has complexity $O(kn^{2})$ for calculating the first $k$ eigenvectors. This holds for Laplacian positional embeddings as well as for spectral features such as Heat Kernel Signature (HKS). Tab. S1 presents experiments evaluating the computational and memory requirements.
@@
Table S1: Metrics for different mesh sizes with respect to the number of vertices, for a 1-layer HodgeFormer (HF) end-to-end architecture (training and inference) as well as a standalone HodgeFormer layer (inference). The model operates on vertices with a latent embedding dimension $d=256$ and hidden MLP dimension of $d_{h}=512$. Measured on an Nvidia RTX 4090 GPU.
Appendix B Results
B.1 Robustness Analysis
To further evaluate the robustness of the HodgeFormer architecture, we perform an analysis, where we evaluate on meshes with added noise, different topology, removed triangles as well as on incomplete meshes. Specifically, we produce variants of the Human dataset’s test set as following:
  • Gaussian Noise: We add Gaussian Noise (GN) $\epsilon$, of several levels $\lambda\in{0.005,0.01,0.02}$ calculated w.r.t. the diagonal of the axis-aligned bounding box of each model, i.e. $\epsilon=\lambda\cdot|BB_{diag}|$.
  • QEM Remeshing: We use the models of the original Human dataset from [25] and remesh them using the Quadratic Error Metric (QEM) to different target face resolution, i.e. $[1000,2000]$.
  • Face Removal: We assign to each face a probability $p$ to be randomly removed, and produce dataset variants with different probabilities in the range $[0.01,0.20]$.
  • Patch removal: We assign to each face a probability $p=0.005$ to be randomly selected and remove a large patch of $k$ neighbors where $k\sim\mathcal{U}(8,15)$.
For all methods, with the exception of QEM Remeshing, we use the test set of the Human dataset. Ground truth for evaluation is defined via nearest-neighbor on the original meshes. We evaluate the pre-trained networks reported in Tab. 5, on the mutated datasets and report the results in Tab. S2. The HodgeFormer architecture retains well its performance under noise addition, remeshing, removed faces and added holes.
HUMAN Test Set Variant Accuracy Acc.Drop
Original 90.3% n/a
Gaussian Noise ($\lambda=0.005$) 88.3% 2.0%
Gaussian Noise ($\lambda=0.010$) 87.3% 3.0%
Gaussian Noise ($\lambda=0.020$) 81.1% 9.2%
QEM Remesh ($1000F$) 87.2% 3.1%
QEM Remesh ($2000F$) 86.2% 4.1%
Face Removal ($p=0.04$) 87.8% 2.5%
Face Removal ($p=0.10$) 85.9% 4.4%
Face Removal ($p=0.20$) 81.6% 8.7%
Patch removal 86.8% 3.5%
Table S2: Robustness evaluation of the HodgeFormer architecture. We report the performance of pre-trained networks on mutated variations of the Human dataset test set. The HodgeFormer architecture retains well its performance under noise addition, remeshing, removed faces and added holes.
Figure S4: Mesh segmentation results on selected models from the mutated Human Body test set. The segmentation results remain reasonable and consistent with the initial predictions of the network.
Figure S5: Mesh segmentation results on selected models from the mutated Human Body test set. The segmentation results remain reasonable and consistent with the initial predictions of the network.
Incomplete Meshes: Additionally, we select two mesh models from the test set of the Human Body Dataset, one with accurate and one with inaccurate segmentation results and gradually remove body parts from them, similar to [10]. Then, we produce segmentations using a pre-trained HodgeFormer network from Tab.5 as showcased in Fig. S5. The segmentation results remain reasonable and consistent with the initial predictions of the network on these mesh models, even on the one with inaccurate segmentation.
B.2 Training on One Sample per Class
We repeat the experiment performed by [30], where a model is trained on the 30-class SHREC11 [20] dataset on splits of only one sample per class (split-1). The model follows the same architecture as Sec.5 with a learning rate of $0.0025$ to compensate for the small dataset size. The results are presented in Tab. S3 along with (split-10) results for comparison.
Method SHREC11 SHREC11
(split-10) (split-1)
SubDivNet [12] 99.5% 36.5%
MWFormer [30] 100.0% 41.7%
HodgeFormer (ours) 98.7% 45.9%
Table S3: Classification performance of different methods on the 30-class SHREC11[20] dataset trained on splits of only 1 sample per class (split-1). Results for (split-10) are also reported for comparison.
B.3 Variations Study on Mixing Strategies
This section examines the effect of mixing HodgeFormer and Transformer layers on segmentation performance. Table S4 summarizes the performance on the COSEG Vases dataset under varying layer compositions. We denote $N_{H}$ as the number of HodgeFormer layers and $R_{H:T}=(N_{H}:N_{T})$ the ratio of the number of HodgeFormer layers to Vanilla Transformer layers. As shown in Table S4, carefully mixing Transformer with HodgeFormer layers has an effect on the final model performance. In general, from the conducted experiments, different mixing strategies would fit different datasets. For the COSEG Vases the best result (93.04%) was obtained by adding two Transformer layers on top of the HodgeFormer layers. Additional layouts of HodgeFormer and Transformer layers are showcased in Sec. B.6.
Layers Vases (COSEG)
@@ B.3 Variations Study on Mixing Strategies
Table S4: Variation experiments comparing different strategies of mixing HodgeFormer with Transformer layers and their effect on the performance, evaluated on the COSEG Vases dataset.
B.4 Variation Study on Label Smoothing
This section examines the effect of label smoothing in the cross-entropy loss on the model’s performance for classification and segmentation tasks. Table S5 reports reports classification results on SHREC11 and segmentation results on COSEG Vases under different label smoothing values.
Label Smoothing SHREC11 Vases
(split-10) (COSEG)
$0.00$ 97.50% 94.03%
$0.05$ 98.33% 94.08%
$0.10$ 98.67% 94.12%
$0.20$ 98.70% 94.30%
$0.40$ 98.33% 93.25%
Table S5: Variation study on the effect of label smoothing on the model’s performance for classification and segmentation tasks.
B.5 Eigenfunctions of the corresponding learned operator
We compute the eigenfunctions of the corresponding learned operators and visualize them on the mesh geometry, providing insights into how the HodgeFormer layer captures and processes geometric features at different scales and orientations. Figure S6 at the top, presents the learned attention maps for two sample meshes of COSEG Aliens dataset measured in layer 6 of the architecture. Figure S6 at the bottom presents the first five eigenfunctions of a COSEG Aliens dataset sample.
Figure S6: Qualitative illustration. Above: Learned attention maps from two samples of the COSEG Aliens dataset. Below: First few eigenfunctions of the corresponding learned operator, $L_{6},head_{1}$ for the 1st sample.
B.6 Hodgeformer & Transformer Layer Layouts
This section includes visualizations of HodgeFormer architecture variations and interleaving strategies, along with notation clarifications. $N_{H}$ corresponds to the number of HodgeFormer transformer layers and $R_{(N_{H},N_{T})}$ corresponds to the ratio of HodgeFormers to vanilla transformers.
@@ B.6 Hodgeformer & Transformer Layer Layouts
Figure S8: HodgeFormer architecture variations. For Human dataset, our highest performing model was a $6$-layered HodgeFormer model mixed with 2 Transformer layers in a $2:1$ ratio, as presented in (a), whereas for Vases, Chairs and Aliens datasets, our highest performing model was a $6$-layered HodgeFormer model followed by two Transformer layers, as presented (b).